OpenStudy (anonymous):

what formula do I use for the problem What is the 24th term of the arithmetic sequence 10, 16, 22, …

OpenStudy (anonymous):

xn+1 = xn+6

OpenStudy (anonymous):

\[x_{n+1}=x_{n}+6\]

OpenStudy (anonymous):

\[x_{n} \epsilon \mathbb{R}\]